OpenStudy (anonymous):

ow do you do (5/4)^-2

OpenStudy (anonymous):

1/ (5/4)^2

OpenStudy (anonymous):

\[(\frac{5}{4})^{-2}=(\frac{4}{5})^2=\frac{16}{25}\]

OpenStudy (anonymous):

Then how do you do (5/4)^2

OpenStudy (anonymous):

negative exponent means take the reciprocal. flip it. square by multiplying by itself.

OpenStudy (anonymous):

Ahah!

OpenStudy (anonymous):

easy to square a fraction

OpenStudy (anonymous):

what about 1/(36^-1/2)

OpenStudy (anonymous):

ok lets go slow. the exponent is negative, so get it out of the denominator and put it in the numerator. then the exponent will be positive.

OpenStudy (anonymous):

so its simply 36^1/2?

OpenStudy (anonymous):

\[\frac{1}{64^{-\frac{1}{2}}}=64^{\frac{1}{2}}\]

OpenStudy (anonymous):

yes. you are right

OpenStudy (anonymous):

i meant 36 not 64!

OpenStudy (anonymous):

Power over root right?

OpenStudy (anonymous):

zactly

OpenStudy (anonymous):

Thanks, I got some more Q's, but ill make a new thread so i can medal more peeps

OpenStudy (anonymous):

Cool!